# MiLiC-Eval: Benchmarking Multilingual LLMs for China’s Minority Languages

### v1

@@ Abstract
Large language models (LLMs) excel in high-resource languages but struggle with low-resource languages (LRLs), particularly those spoken by minority communities in China, such as Tibetan, Uyghur, Kazakh, and Mongolian. To systematically track the progress in these languages, we introduce MiLiC-Eval, a benchmark designed for minority languages in China, featuring 24K instances across 9 tasks. MiLiC-Eval focuses on underrepresented writing systems and provides a fine-grained assessment of linguistic and problem-solving skills. Our evaluation reveals that LLMs perform poorly on syntax-intensive tasks and multi-script languages. We further demonstrate how MiLiC-Eval can help advance LRL research in handling diverse writing systems and understanding the process of language adaptation. [^1] Our data and code are available at https://github.com/luciusssss/MiLiC-Eval.
@@
Chen Zhang,  Mingxu Tao,   Zhiyuan Liao,  Yansong Fengthanks: Corresponding author. Peking University {zhangch,thomastao,fengyansong}@pku.edu.cn liaozy@stu.pku.edu.cn
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in high-resource languages such as English and Chinese, demonstrating the ability to perform sophisticated tasks including creative writing Gómez-Rodríguez and Williams (2023), complex planning Huang et al. (2024), and scientific reasoning Wang et al. (2024c); OpenAI et al. (2024b). However, thousands of low-resource languages (LRLs) remain under-explored by LLMs, which continue to exhibit significant limitations in their basic linguistic capabilities for these languages Alam et al. (2024).
This challenge is particularly pronounced for minority languages spoken in China, such as Tibetan (bo), Uyghur (ug), Kazakh (kk), and Mongolian (mn). Despite being spoken by tens of millions, these languages are largely marginalized in NLP research due to their limited digital representation and the scarcity of training data Zhang et al. (2024b). Furthermore, their use of non-Latin scripts, including non-mainstream ones like those in Mongolian and Kazakh, poses additional challenges for tokenization and language modeling.
@@
Currently, no standardized benchmarks exist for evaluating LLMs on these languages. To fill this gap, we introduce MiLiC-Eval, a comprehensive framework with 24K instances across 9 tasks, focusing on 4 Minority Languages in China. It overcomes limitations in previous LRL benchmarks such as inadequate attention to low-resource writing systems and unsystematic task lineups Hu et al. (2020); Ahuja et al. (2023); Zhang et al. (2024d). MiLiC-Eval offers the following key features: (1) Focus on underrepresented writing systems such as traditional Mongolian and Tibetan scripts; (2) Fine-grained skill evaluation, assessing both linguistic and problem-solving abilities, from vocabulary to symbolic reasoning, as shown in Figure 1; (3) Faithful evaluation with parallel tasks across languages and formats.
Our evaluation of state-of-the-art multilingual LLMs reveals that they especially struggle with languages using less common writing systems such as Mongolian. Skill-wise evaluation shows that LLMs have basic lexical knowledge in these languages and can perform simple tasks such as topic modeling. However, they still lack abilities of generation and translation, which require syntactical knowledge intensively.
@@ 2 MiLiC-Eval
We introduce MiLiC-Eval, the first standardized benchmark designed to evaluate LLMs’ performance on four Minority Languages in China. The benchmark includes 9 distinct task with 24K instances. Basic information about each task is provided in Table 1.
@@ 2.1 Design Principles
While current LLMs have made encouraging progress in supporting LRLs Ji et al. (2024), there has been limited focus on improving their handling of diverse, less common writing systems, particularly those that vary across communities.
The four languages in MiLiC-Eval all use non-Latin scripts, which are poorly supported by English-dominant LLMs. Notably, MiLiC-Eval includes two languages that adopt different writing systems in different communities: The Kazakh and Mongolian communities in China use the Arabic Kazakh script and traditional Mongolian script respectively, instead of the mainstream Cyrillic script. We are the first to benchmark these less-common writing systems. MiLiC-Eval serves as a testbed for evaluating LLMs’ ability to handle less-represented writing systems.
@@
For certain task types, MiLiC-Eval provides multiple subsets for more fine-grained evaluation, such as the separate MT subsets for translating formal texts and colloquial dialogues.
Cross-Language and Cross-Task Parallelism
MiLiC-Eval is highly parallel across both languages and tasks. For six tasks in it, testing instances are provided in Chinese, English, and four minority languages, enabling a clear comparison between high- and low-resource languages. This parallelism also helps researchers from diverse linguistic backgrounds interpret the results.
Additionally, MiLiC-Eval demonstrates task parallelism by using the same text for multiple tasks. This cross-task parallelism enables a more faithful evaluation of LLMs, reducing reliance on shortcuts or superficial patterns. We report the parallelity of tasks in Table 1.
@@ 2.2 Tasks and Dataset Curation
Here we introduce the tasks in MiLiC-Eval and describe the procedure of dataset curation. See details of dataset collection in Appendix B.
@@
Accurately interpreting words is a prerequisite for language comprehension. Though straightforward, vocabulary understanding in MiLiC-Eval specifically evaluates whether LLMs truly grasp the lexicons of LRLs. This task presents multi-choice questions where LLMs must select the correct meaning of a given word from four options. The word-meaning pairs are sourced from Wiktionary [^2] https://en.wiktionary.org/, with distractors sampled from words sharing the same part of speech. We construct 1,000 questions for each language.
@@
Topic classification is a fundamental NLU task that tests whether LLMs can identify the topic of a given text in a minority language. Existing multilingual datasets, like SIB-200 Adelani et al. (2024) and TAXI-1500 Ma et al. (2023), have broad coverage but exclude several languages in our study. While datasets like WCM Yang et al. (2022) and MiTC Deng et al. (2023) exist for these languages, our manual audit reveals issues, such as errors from the automated collection, and indistinguishable labels (e.g., culture vs. literature).
To address these issues, MiLiC-Eval introduces two subsets: Sentence and Passage. The Sentence subset, with 492 instances per language, builds on SIB-200, which includes sentences from Wikipedia, allowing for parallel comparisons across languages. We expand this subset to include Mongolian in the traditional script and Kazakh in the Arabic script  [^3] We transliterate the Cyrillic Kazakh instances in SIB-200 into the Arabic Kazakh script using rules. Since there is no exact mapping for transliteration from Cyrillic to traditional Mongolian, we recruit native speakers from Inner Mongolia to manually transliterate the Cyrillic Mongolian instances in SIB-200 into the traditional Mongolian script.. The Passage subset, with 600 instances per language, evaluates language understanding with passages natively written in the target languages. It is reconstructed from WCM and MiTC by manually removing errors and indistinguishable labels.
@@
Title generation task evaluates LLMs’ ability to generate coherent and concise texts in the target languages after understanding longer documents. Following the approach of XLSUM Hasan et al. (2021), a multilingual summarization dataset, we collect title-article pairs from news websites originally written in minority languages. To avoid contamination from the MC2 corpus Zhang et al. (2024b), which includes news texts until November 2023, we ensure the articles are published after March 2024. In total, we gather 1,000 instances per language.
@@
Machine translation (MT) enables communication across languages, with Flores+ NLLB Team et al. (2024) offering the widest language coverage. However, it lacks Kazakh in the Arabic script and Mongolian in the traditional Mongolian script, and primarily includes formal texts from Wikipedia, overlooking colloquial language use. MiLiC-Eval addresses these gaps with two MT subsets: Article and Dialogue. The Article subset extends Flores+ to include all four languages in our study. We reuse the translated instances in the Topic Classification (Sentence) task as it uses the same set of texts as Flores+. The Dialogue subset assesses MT performance on colloquial texts using dialogues from the reading comprehension task. We split the dialogues into sentences, resulting in 773 instances per language.
@@ 2.3 Evaluation
MiLiC-Eval is a benchmark designed specifically for the evaluation of LLMs. Unlike traditional supervised learning, LLMs require only a few in-context learning (ICL) exemplars to perform tasks effectively. For each task, we provide a small set of instances as ICL exemplars, forming the training set. Three training sets are provided for each task, and we recommend running experiments multiple times with different sets to reduce variance.
@@ 3 Evaluating Multilingual LLMs
We conduct a systematic evaluation of existing multilingual LLMs with MiLiC-Eval and analyze the current progress of China’s minority languages.
@@ 3.1 Experimental Setups
We evaluate two sets of LLMs: (1) native multilingual LLMs, which acquire their multilingual abilities through pretraining, including Llama-3.1 Dubey et al. (2024), Aya-Expanse Dang et al. (2024), Gemma-2 Gemma Team et al. (2024), Qwen-2.5 Yang et al. (2024), Ministral Mistral Team (2024), and GPT-4o-mini OpenAI et al. (2024a); (2) multilingually-adapted LLMs, whose multilingual capabilities are enhanced via multilingual continual pretraining, including EMMA-500 Ji et al. (2024), BayLing-2 Zhang et al. (2024c), LLaMAX-3 Lu et al. (2024), and TowerInstruct Alves et al. (2024). Among these models, only EMMA-500 and BayLing-2 claim using training data for the minority languages of our study.
We mainly evaluate the version with fewer than 10B parameters for each model. Additionally, we evaluate all available versions within the Qwen-2.5 series to investigate the impact of model size. See analyses regarding the effects of model sizes in Appendix D.2.
@@
Following previous works Shi et al. (2023); Asai et al. (2024), we use English as the prompting language. For LLM with fewer than 10B parameters, we run the experiments three times and report the mean results. For LLMs larger than 10B, we only run the experiments once due to limited computational resources. See details in Appendix C.
Figure 2: Skill-wise scores of the top-5 best-performing LLMs on MiLiC-Eval. We report the average scores on the four minority languages.
@@ 3.2 Results
We present the average performance across all tasks for each language in Table 2. Skill-wise performance for the top-5 best-performing LLMs is depicted in Figure 2. For detailed scores, see Table 6 in the Appendix.
@@
Moreover, the performance is severely unbalanced across the four LRLs. As shown in Table 2, current LLMs exhibit a decent understanding of Uyghur, and a preliminary understanding of Tibetan and Kazakh, but struggle with comprehending Mongolian. Even the best-performing model, GPT-4o-mini, only achieves performance slightly better than random on several Mongolian tasks. The low performance of minority languages is partially attributed to their underrepresented writing systems, which we will further discuss in Section 4.1.
@@
As illustrated in Figure 2, LLMs exhibit disparities across various levels of skills in LRLs. Regarding linguistic competence, LLMs demonstrate an ability to comprehend the semantics of input texts by recognizing the meanings of a limited set of words. Furthermore, they show preliminary proficiency in pragmatics and interactive communication. However, their grasp of syntax appears to be shallow, as they struggle to output coherent and grammatically correct sentences in the target languages.
From a problem-solving perspective, current LLMs demonstrate the ability to model the topics of texts in LRLs at a reasonable level. However, their capacity for fine-grained contextual understanding is more limited. Additionally, LLMs face great challenges in both generation and translation tasks. Interestingly, we observe that symbolic reasoning in English can be transferred to LRLs when English is used as the CoT for solving math problems.
@@ 4.1 Underrepresented Writing Systems
We examine tokenization efficiency using parallel sentences from the Machine Translation (Article) task in MiLiC-Eval. We calculate the average number of tokens required to encode sentences in different languages or scripts. As shown in Figure 3, the average token count for English (27 tokens) and Chinese (31 tokens) is much lower compared to the four minority languages, particularly Tibetan and traditional Mongolian, which require 70-430 tokens per sentence. Among the evaluated LLMs, Gemma-2 shows the highest efficiency, which may account for its superior overall performance on MiLiC-Eval. For languages that employ multiple writing systems, tokenization efficiency for less common scripts is often considerably lower than for more widely used scripts. For example, GPT-4o-mini requires 432 tokens for a sentence in traditional Mongolian, eight times the number of tokens needed for the Cyrillic script.
@@
We observe that current LLMs often exhibit code-switching errors when generating content in minority languages, i.e. switching to other languages or scripts during generation. This issue is especially prevalent in Kazakh and Mongolian, both of which employ less commonly used writing systems. For instance, in the title generation task, the GPT-4o-mini model switches to Cyrillic, the more widely used script for both languages, in 36% of Mongolian cases and 95% of Kazakh cases. This problem significantly hurts model performance on generation tasks and may lead to confusion to users when deployed in real-world applications.

### v2

@@ Abstract
Large language models (LLMs) excel in high-resource languages but struggle with low-resource languages (LRLs), particularly those spoken by minority communities in China, such as Tibetan, Uyghur, Kazakh, and Mongolian. To systematically track the progress in these languages, we introduce MiLiC-Eval, a benchmark designed for minority languages in China, featuring 24K instances across 9 tasks. MiLiC-Eval focuses on underrepresented writing systems. Its parallelism between tasks and languages can provide a faithful and fine-grained assessment of linguistic and problem-solving skills. Our evaluation reveals that open-source LLMs perform poorly on syntax-intensive tasks and multi-script languages. We further demonstrate how MiLiC-Eval can help advance LRL research in handling diverse writing systems and understanding the process of language adaptation  [^1] Our data and code are available at https://github.com/luciusssss/MiLiC-Eval..
@@
Chen Zhang,  Mingxu Tao,   Zhiyuan Liao,  Yansong Fengthanks: Corresponding author. Wangxuan Institute of Computer Technology, Peking University {zhangch,thomastao,fengyansong}@pku.edu.cn liaozy@stu.pku.edu.cn
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in high-resource languages such as English and Chinese, demonstrating the ability to perform sophisticated tasks, including creative writing Gómez-Rodríguez and Williams (2023), complex planning Huang et al. (2024), and scientific reasoning Wang et al. (2024c); OpenAI et al. (2024b); Guo et al. (2025). However, thousands of low-resource languages (LRLs) remain under-explored by LLMs, which exhibit significant limitations in their basic linguistic capabilities for these languages Alam et al. (2024).
This challenge is particularly pronounced for minority languages spoken in China, such as Tibetan (bo), Uyghur (ug), Kazakh (kk), and Mongolian (mn). Although spoken by tens of millions, these languages are largely marginalized in NLP research due to their limited digital representation and the scarcity of training data Zhang et al. (2024b). Furthermore, their use of non-Latin scripts, including non-mainstream ones like those in Mongolian and Kazakh, poses additional challenges for tokenization and language modeling.
@@
Currently, there are no standardized benchmarks for evaluating LLMs on these languages. To fill this gap, we introduce MiLiC-Eval, a comprehensive framework with 24K instances across 9 tasks, focusing on 4 Minority Languages in China. It overcomes limitations in previous LRL benchmarks such as inadequate attention to low-resource writing systems and unsystematic task lineups Hu et al. (2020); Ahuja et al. (2023); Zhang et al. (2024d). MiLiC-Eval offers the following key features: (1) Focus on underrepresented writing systems such as traditional Mongolian and Tibetan scripts; (2) Faithful evaluation with parallel tasks across languages and formats; (3) Fine-grained skill evaluation, assessing both linguistic and problem-solving abilities, from vocabulary to symbolic reasoning, as shown in Figure 1.
Our evaluation of open-source multilingual LLMs reveals that they especially struggle with languages using less common writing systems, such as Mongolian. Skill-wise evaluation shows that LLMs have basic lexical knowledge in these languages and can perform simple tasks such as topic modeling. However, they still lack the abilities for generation and translation, which require intensive syntactic knowledge.
@@ 2 MiLiC-Eval
We introduce MiLiC-Eval, the first standardized benchmark for evaluating LLMs’ performance on four minority languages in China. The benchmark includes 9 distinct tasks with 24K instances. Basic information about each task is provided in Table 1.
@@ 2.1 Design Principles
While current LLMs have made encouraging progress in supporting LRLs Ji et al. (2024); Yang et al. (2025), there has been limited focus on improving their handling of diverse, less common writing systems, particularly those that vary across communities.
The four languages in MiLiC-Eval all use non-Latin scripts, which are poorly supported by English-dominant LLMs. Notably, MiLiC-Eval includes two languages that adopt different writing systems in different communities: The Kazakh and Mongolian communities in China use the Arabic Kazakh script and traditional Mongolian script, respectively, instead of the mainstream Cyrillic script. We are the first to benchmark these less-common writing systems. MiLiC-Eval can serve as a testbed for evaluating LLMs’ ability to handle less-represented writing systems.
Cross-Language and Cross-Task Parallelism
MiLiC-Eval is highly parallel across both languages and tasks. For six tasks in MiLiC-Eval, testing instances are provided in Chinese, English, and four minority languages, enabling a clear comparison between high- and low-resource languages. This parallelism also helps researchers from diverse linguistic backgrounds interpret the results.
Additionally, MiLiC-Eval demonstrates task parallelism by using the same text for multiple tasks. This cross-task parallelism enables a more faithful evaluation of LLMs, reducing the reliance on shortcuts or superficial patterns. We report the parallelism of tasks in Table 1.
@@
For certain task types, MiLiC-Eval provides multiple subsets for more fine-grained evaluation, such as the separate machine translation subsets for translating formal texts and colloquial dialogues.
@@ 2.2 Tasks and Dataset Curation
Here we introduce the tasks in MiLiC-Eval and describe the procedure of dataset curation. In Figure 2, we illustrate how each task in MiLiC-Eval is constructed. See details of dataset collection in Appendix B.
Figure 2: The methods of dataset construction for each task in MiLiC-Eval.
@@
Accurately interpreting words is a prerequisite for language comprehension. Although straightforward, vocabulary understanding in MiLiC-Eval specifically evaluates whether LLMs truly grasp the lexicons of LRLs. This task presents multi-choice questions where LLMs must select the correct meaning of a given word from four options. The word-meaning pairs are sourced from Wiktionary [^2] https://en.wiktionary.org/, with distractors sampled from words sharing the same part of speech. We construct 1,000 questions for each language.
@@
Topic classification is a fundamental NLU task that tests whether LLMs can identify the topic of a given text in a minority language. Existing multilingual datasets, such as SIB-200 Adelani et al. (2024) and TAXI-1500 Ma et al. (2023), have broad coverage but exclude several languages in our study. While datasets like WCM Yang et al. (2022) and MiTC Deng et al. (2023) exist for these languages, our manual audit reveals quality issues, such as errors from the automated collection, and indistinguishable labels (e.g., culture vs. literature).
To address these issues, MiLiC-Eval introduces two subsets: Sentence and Passage. The Sentence subset, with 492 instances per language, builds upon SIB-200, which includes sentences from Wikipedia, allowing for parallel comparisons across languages. We expand this subset to include Mongolian in the traditional script and Kazakh in the Arabic script  [^3] We transliterate the Cyrillic Kazakh instances in SIB-200 into the Arabic Kazakh script using rules. Since there is no exact mapping for transliteration from Cyrillic to traditional Mongolian, we recruit native speakers from Inner Mongolia to manually transliterate the Cyrillic Mongolian instances in SIB-200 into the traditional Mongolian script.. The Passage subset, with 600 instances per language, evaluates language understanding with passages originally written in the target languages. It is reconstructed from WCM and MiTC by manually removing errors and indistinguishable labels.
@@
Title generation task evaluates LLMs’ ability to generate coherent and concise texts in the target languages after understanding longer documents. Following the approach of XLSUM Hasan et al. (2021), a multilingual summarization dataset, we collect title-article pairs from news websites originally written in minority languages. To avoid contamination from the MC2 corpus Zhang et al. (2024b), which includes news texts until November 2023, we ensure the articles are published after March 2024. In total, we collect 1,000 instances per language.
@@
Machine translation (MT) enables communication across languages, with Flores+ NLLB Team et al. (2024) offering the widest language coverage. However, it lacks Kazakh in Arabic script and Mongolian in traditional Mongolian script, and mainly includes formal texts from Wikipedia, overlooking colloquial language use. MiLiC-Eval addresses these gaps with two MT subsets: Article and Dialogue. The Article subset extends Flores+ to include all four languages in our study. We reuse the translated instances in the Topic Classification (Sentence) task as it uses the same set of texts as Flores+. The Dialogue subset assesses MT performance on colloquial texts using dialogues from the reading comprehension task. We split the dialogues into sentences, resulting in 773 instances per language.
@@ 2.3 Evaluation
MiLiC-Eval is a benchmark designed specifically for the evaluation of LLMs. Unlike traditional supervised learning, LLMs require only a few in-context learning (ICL) examples to perform tasks effectively. For each task, we provide a small set of instances as ICL examples, forming the training set. Three training sets are provided for each task, and we recommend running experiments multiple times with different sets to reduce variance.
@@
Figure 3: Skill-wise scores of the best-performing LLMs on MiLiC-Eval. We report the average scores on the four minority languages.
@@ 3 Evaluating Multilingual LLMs
We conduct a systematic evaluation of existing multilingual LLMs with MiLiC-Eval and analyze the current progress in China’s minority languages.
@@ 3.1 Experimental Setups
We evaluate a series of competitive proprietary and open-source LLMs. The evaluated proprietary LLMs include GPT-4o-mini OpenAI et al. (2024a), GPT-4.1 OpenAI (2025) and Gemini-2.0-Flash Google (2024). Regarding open-source ones, we evaluate two sets of LLMs: (1) native multilingual LLMs, which acquire their multilingual abilities through pretraining, including Llama-3.1 Dubey et al. (2024), Aya-Expanse Dang et al. (2024), Gemma-2 Gemma Team et al. (2024), Gemma-3 GemmaTeam et al. (2025), Qwen-2.5 Yang et al. (2024), Qwen-3 Yang et al. (2025), and Ministral Mistral Team (2024); (2) multilingually-adapted LLMs, whose multilingual capabilities are enhanced via multilingual continual pretraining, including EMMA-500 Ji et al. (2024), BayLing-2 Zhang et al. (2024c), LLaMAX-3 Lu et al. (2024), and TowerInstruct Alves et al. (2024).
Among these models, only EMMA-500 and BayLing-2 explicitly claim to use training data for the minority languages of our study. We mainly evaluate the version with around 10B parameters for open-source models. Except for EMMA-500, all the evaluated models are instruction-tuned. Additionally, we evaluate all available versions within the Qwen-2.5 series to investigate the impact of model size. See analyses regarding the effects of model sizes in Appendix D.2.
@@
Following previous works Shi et al. (2023); Asai et al. (2024), we use English as the prompting language. For LLMs with fewer than 10B parameters, we run the experiments three times and report the mean results. For LLMs larger than 10B and proprietary LLMs, we only run the experiments once for efficiency. See details in Appendix C.
@@ 3.2 Results
We present the average performance across all tasks for each language in Table 2. Skill-wise performance for the best-performing LLMs is depicted in Figure 3. For detailed scores, see Appendix D.1.
@@
Moreover, the performance is severely unbalanced across the four LRLs. As shown in Table 2, current open-source LLMs exhibit a decent understanding of Uyghur, and a preliminary understanding of Tibetan and Kazakh, but struggle with comprehending Mongolian. Even the best-performing open-source model, Gemma-3, only achieves performance slightly better than random on several Mongolian tasks. The low performance of minority languages is partially attributed to their underrepresented writing systems, which we will discuss further in Section 4.1. Still, it is encouraging to see that proprietary LLMs, especially Gemini-2.0-Flash, show competitive support on the four languages. We hope that this progress could be transferred to open-source models in the future.
@@
As illustrated in Figure 3, most LLMs exhibit disparities across various levels of skills in LRLs. Regarding linguistic competence, LLMs demonstrate an ability to comprehend the semantics of input texts by recognizing the meanings of a limited set of words. Furthermore, they show preliminary proficiency in pragmatics and interactive communication. However, their grasp of syntax appears to be shallow, as they struggle to output coherent and grammatically correct sentences in the target languages.
From a problem-solving perspective, current LLMs demonstrate the ability to model the topics of texts in LRLs at a reasonable level. However, their capacity for fine-grained contextual understanding is still limited. Additionally, LLMs face great challenges in both generation and translation tasks. Regarding symbolic reasoning, LLMs can leverage English CoT to successfully tackle a subset of math problems despite the limited linguistic abilities in these LRLs.
@@ 4.1 Underrepresented Writing Systems
We examine tokenization efficiency using parallel sentences from the Machine Translation (Article) task in MiLiC-Eval. We calculate the average number of tokens required to encode sentences in different languages or scripts. As shown in Figure 4, the average token count for English (27 tokens) and Chinese (31 tokens) is much lower compared to the four minority languages, particularly Tibetan and traditional Mongolian, which require 100-430 tokens per sentence. For languages that employ multiple writing systems, tokenization efficiency for less common scripts is often considerably lower than for more widely used scripts. For example, GPT-4.1 requires 432 tokens for a sentence in traditional Mongolian, eight times the number of tokens needed for the Cyrillic script.
@@
We observe that current LLMs often exhibit code-switching errors when generating content in minority languages, i.e., switching to other languages or scripts during generation, which is also known as language confusion Marchisio et al. (2024). This issue is especially prevalent in the Kazakh and Mongolian languages used in China, both of which employ less commonly used writing systems. For instance, in the title generation task, the GPT-4o-mini model switches to Cyrillic, the more widely used script for both languages, in 36% of Mongolian cases and 95% of Kazakh cases. This problem significantly hurts model performance on generation tasks and may lead to confusion for users when deployed in real-world applications.